a standard deck of 52 cards holds 4 card each of twos through aces in four equal sized suit hearts, diamonds, club, and spades. if you choose a card at random, what is the probability of choosing an ace or a heart?
Answer by stanbon(75887)

a standard deck of 52 cards holds 4 card each of twos through aces in four equal sized suit hearts, diamonds, club, and spades. if you choose a card at random, what is the probability of choosing an ace or a heart?
------
# of ways to succeed = 4 + 13 - 1 = 16
# of possible outcomes:: 52
----
P(are or heart) = 16/52 = 8/26 = 4/13
------------------------------------
Cheers,
Stan H.
-----------------